OpenStudy (anonymous):

Find the domain of the function f( x)= square root x-2+3

OpenStudy (anonymous):

hang on

OpenStudy (e.mccormick):

Well, whatever is under the root must be >=0. So solve for 0, and that gives you what it must be >= to.

OpenStudy (anonymous):

is it square root x-2+3 or just x-2

OpenStudy (e.mccormick):

@pvs285 Good point.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

katlin95 we need that to solve

OpenStudy (anonymous):

square root x-2+3

OpenStudy (anonymous):

is the square root over all of those numbers

OpenStudy (anonymous):

or just over x-2

OpenStudy (anonymous):

yep square root over the whole problem

OpenStudy (anonymous):

x >= -1

OpenStudy (anonymous):

the number is in the top corner of that web sight

jimthompson5910 (jim_thompson5910):

If the equation is \[\Large f(x) = \sqrt{x-2}+3\] then you solve \[\Large x-2 \ge 0\] ------------------------------ If the equation is \[\Large f(x) = \sqrt{x-2+3}\] then you solve \[\Large x-2+3 \ge 0\]

OpenStudy (anonymous):

I know its supposed to be x>-1 but heres my multiple choice that it says

OpenStudy (anonymous):

A. x≥2 B. x<-2 C. x>2 D. x≥2

jimthompson5910 (jim_thompson5910):

solve \[\Large x-2 \ge 0\]

OpenStudy (anonymous):

its x >= 2

OpenStudy (anonymous):

D is supposed to have a -2 sorry lol

OpenStudy (anonymous):

so its A.